# PolicyPad: Collaborative Prototyping of LLM Policies

### v1

@@ 2.4. Tools for LLM Policy Design
Due to the nascency of LLM policy design, there are currently few tools to support policy designers, despite the rising importance of LLM policies (Lam et al., 2024; Huang et al., 2024). Policy Projector by Lam et. al (Lam et al., 2024) supports AI safety practitioners in authoring if-then rules for LLM content moderation by identifying gaps in existing policies through a visual interface. Our work differs in two important ways. First, rather than patching gaps in existing policies, we support the creation of novel policies that can extend beyond what existing policies cover. Second, we specifically support collaborative policy design, which Lam et al. identified as a fruitful area of future work. ConstitutionMaker (Petridis et al., 2024) allows users to turn written critiques of model responses into principles that guide future behavior, but is a tool for LLM personalization rather than policy design. Roleplay-doh (Louie et al., 2024) similarly converts written feedback on LLM behavior into principles, but specifically for domain experts to govern LLM-prompted roleplay. We see this feedback-to-principle interaction as valuable to policy designers as well, and integrate a variant of it into our system.
@@ 3.1. Study Motivation and Procedure
To develop an understanding of real-world LLM policy design practices, we conducted a 15-week-long observational study via contextual inquiry (Beyer and Holtzblatt, 1999) in partnership with a frontier AI lab based in the United States. We wanted to observe how domain experts collaborated to draft domain-specific LLM policies, and any opportunities for improving processes and/or tooling. The lab was convening weekly/twice-a-week virtual workshops (19 workshops total) with 9 experts in clinical mental health (denoted E1–E9) to design new AI policies for model behavior when responding to users’ mental health queries. While all 9 experts were invited to every workshop, there were not 9 attendees every week due to scheduling conflicts. Out of the experts, 6 identified as female and 3 as male. For their highest degrees, 4 held a Ph.D. in clinical psychology, 4 held a Doctor of Clinical Psychology (Psy.D.), and 1 held a Master’s in Clinical Psychology. Experts were all based in the United States. This study was classified as exempt by our institution’s IRB.
@@
At least one member of the research team attended these workshops from January to April 2024. Each workshop was 60–90 minutes in length. One facilitator from either the AI lab or our research team led the workshop with a collaborative policy design activity for the group of experts. These activities revolved around two goals. First, experts developed taxonomies for collections of example mental health-related user queries to an LLM (“scenarios”). Some specific tasks for this goal included deliberating with other experts on taxonomy labels, when to combine and separate labels, and assigning labels to scenarios. Second, experts drafted and voted on desirable rules the model should follow. This included tasks such as reviewing proposed rules in a shared spreadsheet, suggesting modifications, and merging similar rules.
@@
All workshops were recorded and transcribed. As is common in contextual inquiry, team members took notes on observations and asked questions as needed (Beyer and Holtzblatt, 1999). The first author then deductively coded workshop transcripts based on themes identified in our team’s notes, clarifying and iterating on the themes while doing so. We concluded our observational study with a 30-minute semi-structured interview with all the experts in the last workshop. We asked experts to reflect on the workshops and their overall contributions to the policy. The first author used a hybrid inductive-deductive coding process (Fereday and Muir-Cochrane, 2006) to code the interview transcript. This hybrid process allowed us to connect to themes from our workshop data while embracing new themes emerging from the interview. Our final set of themes can be found in Table 2 in Appendix C.
@@ 3.2. Observational Study Results
We used our notes and themes to synthesize four main observations, which we describe in detail below.
@@ 3.2.1. Incomplete feedback loops without model experimentation
Throughout the workshops, experts had visions for how their contributions to the policy—through drafting taxonomies and principles—can impact model behavior. For example, E3 shared that “a clinical minimization [of the user’s feelings] can be helpful, but for the model, that would be hard to decipher” so the group wrote a policy barring the model from engaging in this behavior. However, experts failed to verify whether and how those visions actually came into fruition because they did not have a “policy-informed” model—a model that acts in accordance with the policy—to interact with. This resulted in an incomplete feedback loop. Experts were unable to obtain signals about the effectiveness of their policy contributions and any unintended side effects that may arise, as E9 explains: “Just because you think that might be a good rule, it may have an unanticipated consequence you don’t realize. I think that it would be really helpful to know how these [rules] we’re coming up with actually play out.” E7 agreed and added that direct experimentation with a policy-informed model would allow them to better “see how [a conversational interaction] would play out from the perspective of a user.” Experts had unrestricted internet access throughout the workshops and could test behaviors out on popular chatbots. However, we did not observe instances of this, possibly due to preoccupation with workshop activities, or lack of knowledge about (or in some systems, inability to set) custom system prompts.
@@ 3.2.2. Experts tackled both high-level strategy and low-level semantics
We noticed that experts could easily derail from workflows that would enable them to best contribute their expertise when designing policies. For example, when creating taxonomies for mental health-related user queries, experts spent substantial time wrestling with wording and semantics. Similarly, E9 reflected that much of their time was spent on finding the right wording for taxonomy labels: “we thought needed to not spend forever trying to wordsmith exactly how that needed to appear.” In a separate activity where experts wrote out ideal model responses, E5 agreed that experts should avoid getting stuck in the weeds of low-level wording edits: “It would be more effective at this stage for us to just put our thoughts in about what’s right or wrong, because the time it takes to craft the perfect response is out of scope for this task.” Study facilitators agreed that much of the low-level semantics can be refined post-hoc via LLMs, as long as there are sufficient amounts of expert insight to guide that refinement.
@@ 3.2.3. Scenarios grounded discussions and spurred policy generation
We found that experts engaged in richer discussions that led to insightful policy suggestions after they were given scenarios—examples of user-AI conversations that may arise in real-world use—for reference. For E1, looking at scenarios helped them identify two pieces of information the model should consider in its response: “We need to ask clarifying questions, in particular to clarify the severity and the nature of the dark thoughts this person suggested. Another dimension is to identify how long they’ve been feeling this way and what sources of support they have.” E2 agreed with the need for a severity assessment, suggesting a safety rating scale for the user in case they cannot quickly reach a professional and need an immediate response: “The AI needs to respond, providing resources quickly. Maybe having a rating scale on the scale of zero to 10, how safe are you feeling right now?” Adding on, E3 suggested eliciting the user’s financial ability to pay for therapy and making referrals accordingly: “There might be questions instead like, what is your financial ability to pay for therapy right now? And if it’s within certain ranges, then you might make a community mental health referral, like here’s some Medicare people in your area.” Exploring scenarios helped experts spot recurring problems in model responses and turn them into clear policies. While rules should stay broad enough to be useful, it is unclear how specific they should be. When scenarios show patterns that keep causing issues, they become obvious candidates for new rules. as E7 describes: “I keep seeing this thing over and over and it’s incorrect, so that needs to be a rule.”
@@ 4. LLM Policy Prototyping
Figure 3. Illustration of our envisioned LLM policy prototyping process. Scenarios inform desiderata for the policy via heuristics, which in turn guide the design of the policy. The policy shapes the behavior of a policy-informed LLM, which designers can then test against the scenarios to observe changes in behavior. The process is iterative: feedback from testing may lead to the creation of new scenarios, heuristics, and policy statements.:
@@
Our observations in Section 3 posed challenges that are not foreign to HCI; well-established concepts and methods in UX design and prototyping can offer help in mitigating these challenges. We now use this insight to conceptualize LLM policy prototyping (henceforth “policy prototyping” for brevity), an emerging practice by which groups of individuals can synchronously collaborate on designing an LLM behavioral policy. Specifically, we map observations we identified in Section 3.2 to relevant UX methods, which are then mapped to their usage in policy prototyping. This is depicted in Table 1. We focus our work on low-fidelity policy prototypes (row 2 in Table 1)—artifacts with the primary goal of eliciting and integrating group perspectives on responsible model behavior, rather than a high-fidelity, “production-ready” policy. We leave the translation of low- to high-fidelity policies to future work.
Observation Relevant UX Method UX Definition & Usage Usage in LLM Policy Prototyping Core Literature
Incomplete feedback loops (Section 3.2.1) Rapid Prototyping Tight feedback loops of ideating, implementing, and evaluating design ideas. Allows designers to identify usability issues early, explore alternatives, and align teams to shared visions Tight feedback loops of ideating, drafting, and testing policy statements for quick identification of policy “usability” issues (e.g., unclear statements), characteristics of responsible model behavior, and translations of that behavior into policy. (Lim et al., 2008; Houde and Hill, 1997; Buchenau and Suri, 2000; Camburn et al., 2017; Dow et al., 2010)
Operating at both high and low levels (Section 3.2.2) Low-fidelity prototyping Artifact loosely resembling the final product in terms of look & feel and/or implementation. Cheap to create and iterate upon, ideal for collecting early requirements and feedback. Artifact providing high-level documentation of responsible model behavior to quickly gather and integrate perspectives/feedback. Requires focus on high-level details. We focus on this type of policy prototype in our work. (Rudd et al., 1996; Virzi et al., 1996; Walker et al., 2002)
Operating at both high and low levels (Section 3.2.2) High-fidelity prototyping Artifact closely resembling the final product in terms of look & feel and/or implementation. They are useful for collecting detailed feedback but may be expensive to create. Artifact providing detailed documentation of responsible model behavior to guide alignment efforts, often with polished wording, illustrative examples, legal sign-off, and more. Requires focus on high- and low-level details. Example: OpenAI Model Spec (OpenAI, 2025a). (Rudd et al., 1996; Virzi et al., 1996; Walker et al., 2002)
Scenarios grounded discussions (Section 3.2.3) Storyboarding Concrete representations of users, contexts, and tasks to ground abstract design ideas. Panels add context and illustrate user stories. Promotes reflection and communication among stakeholders. Sample user-AI conversations to ground policy discussions and creation. Conversational turns add context and illustrate sample user & model behaviors. Promotes reflection and communication among stakeholders. (Figma, 2025a; Truong et al., 2006; Hooper and Hsia, 1982; Bodker, 1999; Andersen and Jæger, 1999)
Experts valued synchronous collaboration (Section 3.2.4) Design workshops Common collaborative method for gathering user requirements, studying empirical phenomena, and evaluating interactive systems. Can serve as a field site, research instrument, or a research account. Small-group sessions that serve as a field site for collective ideation and reflection of responsible model behavior in domain-specific use cases. (Rosner et al., 2016; Elsden et al., 2020; Rosenbaum et al., 2002)
Table 1. Mapping of UX methods relevant to insights from our observational study (Section 3) to their usage in LLM policy prototyping.:
@@ 5.1. Iterative Co-Design Sessions with Experts
We provide detailed documentation of how we integrated participants’ feedback to iterate from a basic first version to our final design in Appendix D. In the following section, we illustrate PolicyPad’s capabilities using a system walkthrough.
@@ 5.2. System Walkthrough
PolicyPad can be used by any individual or group who wishes to facilitate a policy prototyping session—whether it be an AI lab, academic group, non-profit, or another organization. As a note on this section’s terminology, we distinguish “facilitators” (those running the policy prototyping session) from “users” (policy designers participating in the session).
Figure 4. Main components of the PolicyPad system. Users can keep track of their policy version in the left sidebar (A) as they collaborative edit the policy in the editor (B). Users can access scenarios via the scenario gallery (C). When they click into a scenario, they can view its full details and explore how the policy-informed model will behave on it via the scenario sidebar (D).:
@@ 5.2.2. Scenario sidebar
A user can browse the scenarios in the gallery and open a scenario in a detailed view (Fig. 4 D). The scenario expands to fill the sidebar with the full user-AI conversation, as well as a brief, AI-generated summary of the conversation’s contents thus far. As the group reads the scenarios, they start to develop ideas for what to include in the policy.
Formally, a scenario in PolicyPad comprises of three parts: the background (all messages in the conversation up until the most recent turn), the newest user message, and the newest AI message. Given the background and the newest user message, the policy-informed model generates the newest AI message.

### v3

@@ 2.4. Tools for LLM Policy Design
Due to the nascency of LLM policy design, there are currently few tools to support policy designers, despite the rising importance of LLM policies (Lam et al., 2024; Huang et al., 2024). Policy Projector by Lam et al. (Lam et al., 2024) supports AI safety practitioners in authoring if-then rules for LLM content moderation. Our work differs in that we specifically support collaborative policy design, which Lam et al. identified as a fruitful area of future work. ConstitutionMaker (Petridis et al., 2024) allows users to turn written critiques of model responses into principles that guide future behavior, but is a tool for LLM personalization rather than policy design. Roleplay-doh (Louie et al., 2024) similarly converts written feedback on LLM behavior into principles, but specifically for domain experts to govern LLM-prompted roleplay. We see this feedback-to-principle interaction as valuable to policy designers as well, and integrate a version of it into our system.
@@ 3.1. Study Motivation and Procedure
To develop an understanding of real-world LLM policy design practices, we conducted a 15-week-long observational study via contextual inquiry (Beyer and Holtzblatt, 1999) in partnership with OpenAI. We wanted to observe how domain experts collaborated to draft domain-specific LLM policies, [^3] Recall that an LLM policy is a set of rules, guidelines, and desiderata that shape model behavior. and any opportunities for improving processes and/or tooling.
We acknowledge that “expertise” is a contested topic, especially in democratic decision-making contexts (Díaz and Goñi, 2024). We use a narrow definition of expert and recruit participants who have 1) at least two years of practical experience conducting client-facing work in their domain, and 2) are pursuing or have completed an advanced degree (Master’s or PhD or equivalent). This combination ensures that policy design is guided by both participants’ real-world experiences and aspects of their formal training.
3.1.1. Background and participants
OpenAI, in collaboration with a small group of external researchers including us, was organizing weekly/twice-a-week virtual workshops (19 workshops total) with 9 experts in clinical mental health (denoted E1–E9) to design new AI policies for model behavior when responding to users’ mental health queries. While all 9 experts were invited to every workshop, there were not 9 attendees every week due to scheduling conflicts. Out of the experts, 6 identified as female and 3 as male. For their highest degrees, 4 held a Ph.D. in clinical psychology, 4 held a Doctor of Clinical Psychology (Psy.D.), and 1 held a Master’s in Clinical Psychology. Experts were all based in the United States. Participant recruitment and payment were handled by OpenAI and its data partners. Data from the workshops were shared among all collaborators, but our data analysis procedures and research outputs were distinct from those of the company. This study was classified as exempt by the University of Washington IRB.
3.1.2. Procedure
@@
At least one member of our research team attended these workshops from January to April 2024. Each workshop was 60–90 minutes in length. One facilitator from either the AI lab or our research team led the workshop with a collaborative policy design activity for the group of experts. These activities revolved around two goals. First, experts developed taxonomies for collections of example mental health-related user queries to an LLM (“scenarios”). Some specific tasks for this goal included deliberating with other experts on taxonomy labels, when to combine and separate labels, and assigning labels to scenarios. Second, experts drafted and voted on desirable rules the model should follow. This included tasks such as reviewing proposed rules in a shared spreadsheet, suggesting modifications, and merging similar rules. The agenda for each workshop was set by either OpenAI researchers or one of its collaborators. Facilitators’ guides for the workshops we organized can be found in Supplementary Materials.
Experts generally tackled the first goal in earlier weeks and the second in later weeks. There was, however, fluid movement and substantial iteration across the two goals, such that some workshops contained activities pertaining to both goals. As the weeks passed, we also moved towards co-designing tooling [^4] Note that this specifically refers to co-designing tooling for collaborative policy design. It is in service of our broader goal of co-designing policies with experts. with experts as we developed a better understanding of the problem space and experts’ pain points. As part of the final three workshops, we reserved some time to show experts initial prototypes of an interactive system to assist with collaborative LLM policy design, collected their feedback, and iterated on the prototype for the following week (see Appendix B for prototype iterations). In the final workshop, we conducted a 30-minute semi-structured exit interview with all experts, asking them to reflect on the workshops and their policy contributions. Figure 3 shows an overview of activities across the 15-week period.
Figure 3. Timeline of activities in our 15-week observational study. During taxonomy development, experts organized and taxonomized a collection of diverse LLM usage scenarios. During rule writing, experts drafted, discussed, and refined rules to govern LLM behavior. During co-design, experts interacted with and gave feedback on prototypes we built of a tool for collaborative policy design. There was fluid movement between taxonomy development and rule-writing, such that some sessions included activities pertaining to both goals.
Weeks represented by numbers 1 to 15 are listed out horizontally. Highlighting underneath shows that taxonomy development spanned weeks 1–6 and 8–10. Rule writing spanned weeks 4–7 and 9–15. Co-design spanned 13–15.
3.1.3. Data analysis
@@
All workshops were recorded and transcribed. As is common in contextual inquiry, team members took observational notes and asked questions as needed (Beyer and Holtzblatt, 1999). The first author then deductively coded workshop transcripts based on themes identified in our team’s notes, clarifying and iterating on the themes while doing so. The first author used a hybrid inductive-deductive process (Fereday and Muir-Cochrane, 2006) to code the transcript of the exit interview. This hybrid process allowed us to connect to themes from our workshop data while embracing new themes emerging from the interview. Our final set of themes can be found in Table 1 in Appendix A.
@@ 3.2. Observational Study Results
We used our notes and themes from all workshops, including our exit interview, to synthesize four main observations, which we describe in detail below. Henceforth, we use the term “policy-informed model” to refer to an LLM that has been instructed to act in accordance to the policy drafted by experts.
@@ 3.2.1. Incomplete feedback loops without model experimentation
Throughout the workshops, experts had visions for how their contributions to the policy through drafting taxonomies and principles can impact model behavior. For example, E3 shared that “a clinical minimization [of the user’s feelings] can be helpful, but for the model, that would be hard to decipher” so the group wrote a policy barring the model from engaging in this behavior. However, they failed to verify whether and how those visions actually came into fruition because they did not have a policy-informed model to interact with. This resulted in an incomplete feedback loop. Experts were unable to obtain signals about the effectiveness of their policy contributions and any unintended side effects that may arise, as E9 explains: “Just because you think that might be a good rule, it may have an unanticipated consequence you don’t realize. I think that it would be really helpful to know how these [rules] we’re coming up with actually play out.” E7 agreed and added that direct experimentation with a policy-informed model would allow them to better “see how [a conversational interaction] would play out from the perspective of a user.” Experts had unrestricted internet access throughout the workshops and could test behaviors out on popular chatbots. However, we did not observe instances of this, possibly due to preoccupation with workshop activities, or lack of knowledge about (or in some systems, inability to set) custom system prompts.
@@ 3.2.2. Experts tackled both high-level strategy and low-level semantics
We noticed that experts could easily derail from workflows that would enable them to best contribute their expertise when designing policies. For example, when creating taxonomies for mental health-related user queries, experts spent substantial time wrestling with wording and semantics. Similarly, E9 reflected that much of their time was spent on finding the right wording for taxonomy labels: “we thought needed to not spend forever trying to wordsmith exactly how that needed to appear.” In a separate activity where experts wrote out ideal model responses, E5 agreed that experts should avoid getting stuck in the weeds of low-level wording edits: “It would be more effective at this stage for us to just put our thoughts in about what’s right or wrong, because the time it takes to craft the perfect response is out of scope for this task.” While important, study facilitators agreed that much of the low-level semantics of the policy can be refined post-hoc via LLMs, as long as there are sufficient amounts of expert insight to guide that refinement.
@@ 3.2.3. Scenarios grounded discussions and spurred policy generation
We found that experts engaged in richer discussions that led to insightful policy suggestions after they were given scenarios, or examples of user-AI conversations that may arise in real-world use, for reference. For E1, looking at scenarios helped them identify two pieces of information the model should consider in its response: “We need to ask clarifying questions, in particular to clarify the severity and the nature of the dark thoughts this person suggested. Another dimension is to identify how long they’ve been feeling this way and what sources of support they have.” E2 agreed with the need for a severity assessment, suggesting a safety rating scale for the user in case they cannot quickly reach a professional and need an immediate response: “The AI needs to respond, providing resources quickly. Maybe having a rating scale on the scale of zero to 10, how safe are you feeling right now?” Adding on, E3 suggested eliciting the user’s financial ability to pay for therapy and making referrals accordingly: “There might be questions instead like, what is your financial ability to pay for therapy right now? And if it’s within certain ranges, then you might make a community mental health referral, like here’s some Medicare people in your area.” Exploring scenarios helped experts spot recurring problems in model responses and turn them into clear policies. While rules should stay broad enough to be useful, it is unclear how specific they should be. When scenarios show patterns that keep causing issues, they become obvious candidates for new rules, as E7 describes: “I keep seeing this thing over and over and it’s incorrect, so that needs to be a rule.”
@@ 4. LLM Policy Prototyping
Our observations in Section 3 posed challenges that are not foreign to HCI; well-established concepts and methods in UX design and prototyping can offer help in mitigating these challenges. Indeed, scholars have been increasingly interested in applying prototyping principles to designing bills and other public policies (Kimbell and Bailey, 2017; Hagan, 2021; Project Let’s Talk Privacy, 2020; Quicksey and Meierling, 2022; Kontschieder, 2018), but minimal efforts have been made to expand these applications to beyond policies for governments. Additionally, while literature in end-user model auditing and red-teaming has allowed non-AI experts to identify problematic model behaviors (see Section 2.3), few methods exist for directing outcomes of efforts to meet increasing demand for iteratively improving LLM policy design (He et al., 2025; Zhang et al., 2025b; Huang et al., 2024; OpenAI, 2025a). Our work seeks to fill this need through co-designing policies with experts.
We now describe LLM policy prototyping (henceforth “policy prototyping” for brevity), an emerging practice by which groups of individuals can synchronously collaborate on designing an LLM behavioral policy. Specifically, we map observations we identified in Section 3.2 to relevant UX methods, which are then mapped to their usage in policy prototyping. For example, to address Section 3.2.1, enabling tight feedback loops between ideation, design, and testing allows policy designers to quickly identify “usability” issues like unclear policy statements, explore policy designs that more effectively achieves desired model behavior, and work collaboratively to address issues as they surface. For Section 3.2.2, experts’ efforts can be better focused by conducting heuristic evaluation on a policy using heuristics that direct experts’ attention to higher-level desiderata (e.g., does the policy draw from real-world practices in their field?) rather than low-level wording edits. Promising uses of scenarios in Section 3.2.3 can be further scaffolded with techniques inspired by storyboarding, where each conversational turn in the scenario grounds policy design ideas in concrete representations of users, contexts, and tasks. Our full mapping is depicted in Table 2 in Appendix C.
@@
In this work, we focus on low-fidelity policy prototypes—artifacts with the primary goal of eliciting and integrating group perspectives on responsible model behavior, rather than a high-fidelity, “production-ready” policy. We leave the translation of low- to high-fidelity policies to future work.
@@
Figure 4. Illustration of our envisioned LLM policy prototyping process. Scenarios inform desiderata for the policy via heuristics, which in turn guide the design of the policy. The policy shapes the behavior of a policy-informed LLM, which designers can then test against the scenarios to observe changes in behavior. The process is iterative: feedback from testing may lead to the creation of new scenarios, heuristics, and policy statements.
A figure illustrating the flow of LLM policy prototyping: Scenarios (1) inform desiderata for the policy via heuristics (2), which in turn guide the design of the policy (3). The policy shapes the behavior of a policy-informed LLM (4), which designers can then test against the scenarios to observe changes in behavior. The process is iterative: feedback from testing may lead to the creation of new scenarios, heuristics, and policy statements.
@@ 5.1. Iterative Co-Design Sessions with Experts
Our first prototype consisted of a simple collaborative editor for policy authoring, along with a sidebar that allowed experts to engage in conversation with the policy-informed model. Experts’ appreciation of the realtime collaborative features and desires for more structured and systematic ways to interact with scenarios led us to the next iteration. Our second prototype introduced a more sophisticated side panel that allowed users to browse scenarios and use them to test model behavior. While experts agreed that this was a significant improvement, they desired closer integration between policy editing and scenario exploration. In our third prototype, we integrated scenarios into the editor as interactive widgets, while keeping the sidebar as a place to view scenario-specific information. Experts appreciated this and suggested ways to flag problematic model responses from the sidebar, as well as cleaning up unnecessary elements from the sidebar. We incorporated these suggestions into our final system. Detailed documentation on how we integrated participants’ feedback across versions, along with prototype screenshots, can be found in Appendix B. In the following section, we illustrate PolicyPad’s capabilities using a system walkthrough.
@@ 5.2. System Walkthrough
PolicyPad can be used by any individual or group who wishes to facilitate a policy prototyping session—whether it be an AI lab, academic group, non-profit, or another organization. To do so, we use a running example of hypothetical LLM policy for financial advice. As a note on this section’s terminology, we distinguish “facilitators” (those running the policy prototyping session) from “users” (policy designers participating in the session).
Figure 5. Main components of the PolicyPad system. Users can keep track of their policy version in the left sidebar (A) as they collaborative edit the policy in the editor (B). Users can access scenarios via the scenario gallery (C). When they click into a scenario, they can view its full details and explore how the policy-informed model will behave on it using the scenario sidebar (D).
Main components of the PolicyPad system. Users can keep track of their policy version in the left sidebar (A) as they collaborative edit the policy in the editor (B). Users can access scenarios via the scenario gallery (C). When they click into a scenario, they can view its full details and explore how the policy-informed model will behave on it using the scenario sidebar (D).
@@ 5.2.2. Scenario sidebar
A user can browse the scenarios in the gallery and open a scenario in a detailed view (Fig. 5 D). The scenario expands to fill the sidebar with the full user-AI conversation, as well as a brief, AI-generated summary [^6] Generated with GPT-4o. More technical details are in Section 5.3. of the conversation’s contents thus far. As the group reads the scenarios, they start to develop ideas for what to include in the policy.
Formally, a scenario in PolicyPad comprises three parts: the background (all messages in the conversation up until the most recent turn), the newest user message, and the newest AI message. Given the background and the newest user message, the policy-informed model [^7] The policy-informed model is an instance of Llama 3.3 70B Instruct. More technical details are in Section 5.3. generates the newest AI message.